The velocity with which body is to be projected vertically up so that distance covered in 5th secondis twice that covered in 6th second is

Grade:11

2 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Chakrala,

 

distance travelled in nth second  = u +a(2n-1)/2

so, S5  = u + 5 *9

S6 = u + 5 * 11

 

also S5 = 2 S6  use these eq to get u.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay

 

 

chakrala mounica
13 years ago

Sir, thank you for sending answers for two questions, i approve with your answers please send me the remaining answer.These answers help me very much,so please send me.